# Electricity Price Forecasting: Bridging Linear Models, Neural Networks and Online Learning

### v1

@@ 5 Results and Discussion
During the first hour of the day, the LEAR has the lowest RMSE. However, its performance worsened afterwards; it even has the worst performance among all hours, starting from hour 15. The second benchmark model, which is DNN, performs poorly in most hours of the day. Except for the hour zero, MLP with Reduced Linear and MLP with Reduced Linear (BOA) generate the most accurate forecast compared to other models; they interchangeably have the lowest RMSE. For instance, MLP with Reduced Linear is the best performing model in hours 5 and 22, with values of 11.66 and 12.46, respectively, while MLP with Reduced Linear (BOA) has the lowest RMSE in hours 2 and the last hour.
@@
It is apparent that the model performance changes from one hour to another. Hence, to select the best performing model, three overall accuracy measures are calculated and the results are presented in Table 1. The green cells in Table 1 correspond to the best performing model for the respective metric, while the red cells depict the worst performing ones.
Table 1: Comparative Forecast Accuracy in terms of RMSE (EUR/MWh), MAE (EUR/MWh) and rMAE of Competing Models
@@
For both the German-Luxembourg and Spanish bidding zones, MLP with Reduced Linear (BOA) is the best performing model according to MAE and rMAE. On the other hand, when considering RMSE, the MLP with Reduced Linear model is the most accurate in the German–Luxembourg market, whereas the emphMLP with Full Linear model achieves the best performance in the Spanish market. It is apparent that all the developed models in this paper outperformed the state-of-the-art benchmark models in both markets.
@@
The RMSE is reduced by 12-13% and the MAE by 15-18% when comparing the best models to the best-performing benchmark, DNN, in the respective market. Additionally, one can notice an overall decrease in both RMSE and MAE in the Spanish market compared to the same models in the German-Luxembourg market. This might be the result of the share of renewables in the electricity generation. [rintamaki2017does] confirmed that wind-powered generation increases the price volatility in Germany. On the other hand, [pereira2017effect] concluded that while wind increases volatility in Spain, the integration of hydro power plants decreases it. Finally, it is worth emphasizing that for the three measures, we can see a significant improvement from the MLP model to the MLP with Reduced Linear. For instance, the RMSE of MLP with Reduced Linear in the Spanish bidding zone is smaller than that of the MLP by 10%. This highlights the importance of the novel architecture that we proposed in improving accuracy compared to a neural network without a skip connection.
@@
Figure 6 presents the heatmap of the p-value of the Diebold Mariano test for the German-Luxembourg bidding zone. The horizontal model represents model A, while the vertical one represents model B. With a 5% significance level, the null hypothesis is rejected if the p-value is less than 0.05. Based on the legend, shades of green indicate rejection of the null, whereas shades of red indicate failure to reject.
@@
Figure 6: Pairwise DM Test P-Values Across Forecasting Models for German-Luxembourg market
MLP with Reduced Linear (BOA) is statistically more accurate than any other model except BOA all. While MLP with Reduced Linear is the best model according to RMSE, DM results show that this model is not statistically better than MLP with Full Linear, MLP with Reduced Linear (BOA), and BOA all. This may be attributed to the choice of loss function in the DM test. Using a squared loss function might change the results. The red color in LEAR and DNN cells confirms the previous remark about them as the worst-performing models, given that DNN is only statically more accurate than LEAR, and LEAR is not statistically better than any model. Similar results are observed for the Spanish market, where the benchmark models are not statistically more precise than any of the developed models.
Figure 7 illustrates the trade-off between the run time of a model and its accuracy. The best performing model in terms of time is a Full Linear (Reduced Linear for Spain) with a running time of 14 seconds for the whole 730 iterations. This result is consistent with expectations, as linear models typically require the shortest execution time ([a]). MLP with Reduced Linear (BOA) is the slowest model within the Pareto frontier models for both markets. At the same time, it is the model with the best MAE among all models.
While the benchmark models are the worst-performing models in terms of accuracy, their performance in terms of time is much worse. It took LEAR about two hours to train and generate forecasts for the two years of testing, whereas DNN needed about five hours in the German-Luxembourg market and more than three hours for Spain.
@@
Figure 7: Runtine and MAE of considered forecasting models with efficient Pareto Frontier.
@@ 6 Conclusion
We adopt an ensemble method, following the fully adaptive BOA approach, to aggregate individual forecast models and produce more accurate predictions. Moreover, a novel nonparametric partial online learning method is introduced to reduce the computational cost, such that the optimal weights of the previous iteration are employed as initialization of the current iteration.
The models’ performance is evaluated against two state-of-the-art benchmarks: LEAR and DNN models, using German-Luxembourg and Spanish bidding zones data, which are among the largest European markets. The measures of accuracy demonstrate the superiority of all the proposed models over the benchmark models with a 12% reduction in RMSE and 18% decrease in MAE for the German-Luxembourg market and a decline of 13% in RMSE and 15% in MAE for Spain. The results of the Diebold Mariano test are consistent with measures of accuracy metrics. Both LEAR and DNN fail to outperform any of the proposed models.
In addition to their strong predictive accuracy, the suggested models achieve an outstanding runtime. It took the fast models only a few seconds to fit and generate the forecast, while even the slowest and most accurate models complete training and predictions within a few minutes. By contrast, the runtimes of benchmark models are substantially higher, such that each one of them took a couple of hours. This suggests that the proposed online learning approach was very effective in reducing computational time.
@@ Acknowledgements
This work was supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) project numbers 520388526 to BE and FZ.
@@
Figure 11: Pairwise DM Test P-Values Across Forecasting Models for Spanish market

### v2

@@ 5.1 Forecasting Accuracy
During the first hour of the day, the LEAR has the lowest RMSE. However, its performance worsened afterwards. The second benchmark model, which is DNN, performs poorly in most hours of the day. Except for the hour zero, MLP with RLin (BOA) generates the most accurate forecast compared to other models as it has the lowest RMSE.
@@
It is apparent that the model performance changes from one hour to another. Hence, to select the best performing model, three overall accuracy measures are calculated and the results are presented in Table 1.
Germany Spain
RMSE MAE rMAE RMSE MAE rMAE
RLin 24.94 13.86 38.01% 16.59 11.97 38.65%
MLP 23.66 13.07 35.82% 15.76 11.57 37.36%
MLP with RLin 23.37 12.39 33.96% 15.94 11.69 37.76%
FLin 24.81 13.68 37.50% 16.56 12.16 39.26%
MLP with FLin 23.12 12.33 33.79% 15.84 11.63 37.57%
RLin with OLS 25.20 13.99 38.34% 16.35 11.81 38.16%
MLP with RLin and OLS 23.08 12.36 33.88% 16.87 12.23 39.51%
RLin (BOA) 23.58 12.65 34.69% 15.83 11.32 36.55%
MLP (BOA) 22.21 11.66 31.97% 14.96 10.90 35.19%
MLP with RLin (BOA) 22.54 11.67 31.98% 14.85 10.80 34.88%
FLin (BOA) 23.37 12.65 34.67% 15.50 11.27 36.41%
MLP with FLin (BOA) 23.15 11.86 32.51% 14.87 10.81 34.90%
RLin with OLS (BOA) 23.66 12.75 34.95% 17.13 11.43 36.91%
MLP with RLin and OLS (BOA) 22.63 11.56 31.68% 15.03 10.82 34.96%
BOA all 22.62 11.70 32.06% 14.93 10.80 34.90%
LEAR 26.57 16.16 44.30% 17.77 13.24 42.75%
DNN 25.05 13.99 38.37% 16.96 12.61 40.71%
Table 1: Comparative Forecast Accuracy in terms of RMSE (EUR/MWh), MAE (EUR/MWh) and rMAE of Competing Models. For each metric, the cell colors range from red (indicating poorer performance) to blue (indicating better performance). The best performing models are highlighted in bold.
@@
For both the German-Luxembourg and Spanish bidding zones, ensemble models generally outperform individual and benchmark models. For instance, in the Spanish market, MLP with RLin (BOA) achieves the lowest RMSE, MAE, and rMAE among all evaluated models. Overall, most of the models developed in this paper outperform the state-of-the-art benchmark models in both markets. In the German market, even a linear model such as FLin demonstrates better forecasting performance than the nonlinear benchmark DNN.
The RMSE is reduced by 11-12% and the MAE by 14-17% when comparing the best model to DNN in the respective market. RMSE and MAE are also reduced when comparing BOA models to their individual counterparts. For instance, there is a 4% to 11% reduction in MAE in the German Luxembourg market. The gains from ensembling vary across model classes and may be attributed to differences in the correlation between forecast errors, as models with different structures usually benefit more from ensembling. For example, the empirical results of [marcjasz2018selection] and [hubicka2018note] suggest that combining models trained on short and long training windows improves accuracy.
@@
Additionally, one can notice an overall decrease in both RMSE and MAE in the Spanish market compared to the same models in the German-Luxembourg market. This might be the result of the share of renewables in the electricity generation. [rintamaki2017does] confirmed that wind-powered generation increases the price volatility in Germany. On the other hand, [pereira2017effect] concluded that while wind increases volatility in Spain, the integration of hydro power plants decreases it.
@@
Figure 6 presents the heatmap of the p-value of the Diebold Mariano test for the German-Luxembourg bidding zone (see Figure 12 in the Appendix for Spain) under the absolute loss. The horizontal model represents model A, while the vertical one represents model B. With a 5% significance level, the null hypothesis is rejected if the p-value is less than 0.05. Based on the legend, shades of green indicate rejection of the null, whereas shades of red indicate failure to reject.
@@
Figure 6: Pairwise DM Test P-Values Across Forecasting Models for German-Luxembourg market using L1 loss
MLP with RLin and OLS (BOA) is more accurate than any other model except MLP (BOA), MLP with RLin (BOA) , and BOA all , with the difference being statistically significant. On the other hand, the benchmark models perform poorly. The green vertical cells for LEAR and DNN in Figure 6 imply most of the proposed models generates statistically significantly more accurate forecasts than the benchmark models, confirming the previous observation that these models are the worst-performing ones. Moreover, the horizontal cells for the benchmark models indicate that DNN is only statistically significantly more precise than LEAR, whereas LEAR is not statistically significantly more accurate than any other model.
Since the loss function is a key component of the DM test, we implement the DM test under squared loss to examine whether the conclusions about predictive ability change when changing losses. The DM output under the squared loss for different markets are dipicted in Figure 13 and 14 in the Appendix, where the majority of the developed models exhibit statistically significant improvements in forecast accuracy compared to the benchmark models. This confirms that the superior forecasting performance of our models relative to the benchmarks is robust across different markets and choices of loss functions.
5.2 Computational Cost vs Accuracy
Figure 7 illustrates the trade-off between the runtime of a model and its accuracy, where runtime denotes the time required to execute the full forecasting study on the test dataset [^2] The recorded runtime of the proposed and benchmark models was measured on a MacBook Pro (2020) equipped with an Apple M1 processor (8 CPU cores) and 16 GB of unified memory.. The execution time of the proposed models ranges from 8 seconds to about 5 minutes in the German–Luxembourg market and from 7 seconds to 3 minutes and half in the Spanish market. The best performing models in terms of time are generally linear models. In contrast, the ensemble models are the slowest with the highest forecast accuracy.
@@ 5.2 Computational Cost vs Accuracy
Figure 7: Runtime and MAE of the considered forecasting models with the efficient Pareto frontier. The red markers correspond to Pareto-optimal models, while the blue markers represent non-Pareto models.
While the benchmark models are the worst-performing models in terms of accuracy, their performance in terms of time is much worse. It took LEAR about 100 to 108 minutes to train and generate forecasts for the two years of testing, whereas DNN needed about five hours in the German-Luxembourg market and about 8 hours for Spain.
5.3 Impact of Study Design and Hyperparameters
This section examines the impact of the online approach implementation and evaluates the model’s robustness to variations in hyperparameters and time periods. All the analyses are conducted using one of the best-performing specifications, namely MLP with RLin model. In the time-period analysis, we additionally consider the MLP with RLin (BOA). All experiments are carried out using the German–Luxembourg dataset.
To study the impact of online learning on accuracy and computational time, we train MLP with RLin in three different settings without implementing online learning. Particularly, no warm start is applied and a fixed set of hyperparameters is used for all training windows (initial window and updated windows).
In the first setting, MLP with RLin is trained using the hyperparameters of the initial window of the online learning, i.e., the number of epochs set to 60, and the remaining hyperparameters are tuned based on the range specified in the first column of Table 4 in the Appendix. This model is referred to as MLP with RLin (Initial). The second setting is where the model is trained using the hyperparameter of updated windows, that is, the number of epochs is set to 10 and the other hyperparameters are tuned according to the Rangeup column in Table 4. This model is denoted as MLP with RLin (Update). The last setting is similar to the first one, with the difference that the number of epochs is treated as a tunable hyperparameter, with a range from 5 to 75. This model is named MLP with RLin (Tuned).
The above three discussed models without online learning are compared with the original model that employs online learning, hereafter denoted as MLP with RLin. The corresponding results are reported in Table 2.
Model MAE RMSE Runtime (sec)
MLP with RLin 12.39 23.37 39.60
MLP with RLin (Tuned) 12.78 23.72 276.70
MLP with RLin (Update) 12.81 23.90 209.60
MLP with RLin (Initial) 14.05 26.69 17.50
Table 2: Performance Comparison With and Without Online Learning. Runtime represents the time (in seconds) required to complete the full forecasting study across the 730 training windows in the test dataset. For each column the cell colors range from red (indicating poorer performance) to blue (indicating better performance). The best performing models are highlighted in bold.
MLP with RLin (Initial) is the best performing model in terms of computational time, while it has the worst performance in terms of accuracy. MLP with RLin (update) and MLP with RLin (tuned) are much better in terms of forecast performance, but they require a long time to train the model. On the other hand, MLP with RLin, which represents the model with the online approach, is very fast and yields the most accurate forecast. This outcome is particularly noteworthy, since [ash2020warm] argues that warm-start strategies reduce computation time and negatively affect generalization performance, whereas the present results demonstrate that combining warm-starting with window-specific adaptation produces simultaneous gains in accuracy and computational efficiency.
To investigate the impact of hyperparameters on model performance, the following analysis is conducted. The optimal hyperparameters selected based on the validation dataset are fixed, while the hyperparameter under study is allowed to vary. Each time the model is trained with a different value of the free hyperparameter from a predefined grid to examine how forecasting accuracy and computational time change across different configurations. Figure 8 displays the results of the sensitivity analysis for the four main hyperparameters, examined separately.
Figure 8: Sensitivity analysis of the main hyperparameters. Blue markers indicate Pareto-optimal configurations, whereas gold markers denote dominated ones.
The left two panels show the results for window size and number of epochs for the initial training window. These two hyperparameters seem to have a limited impact on performance, as the time and accuracy do not vary significantly when the hyperparameters change across the considered values. On the other hand, the window size and number of epochs for the updated training window (right panels) affect the model performance substantially. For instance, when the window size is 1 (corresponding to the case of pure online learning), the MAE is 27 with a running time of less than 4 seconds only. Later, one can see a clear trend where the increase in window size is associated with a decrease in MAE and an increase in computation time. A similar trend can be observed in the number of epochs for the updated training window.
After investigating the effect of online learning and hyperparameters on the performance of the model, we now assess the robustness of the model across different time periods. Throughout this work, the analysis of all model are carried out using a validation and test dataset specified in Subsection 4.2. In the present analysis, however, we modify the data split: the validation dataset covers the period from 2020-01-17 to 2021-01-15, while the test dataset spans the subsequent three years, from 2021-01-16 to 2024-01-15. Train data, on the other hand, are the observations that precede the target observation to be forecasted, following the online scheme framework adopted throughout the study. The importance of the new split is that the test dataset covers one of the largest European energy crises in recent years. This setup, therefore, provides a rigorous evaluation of the model’s robustness under extreme market conditions.
Using the new validation dataset, the optimal hyperparameters are selected based on the same ranges specified in Table 4 in the Appendix, with the exception of $D_{\text{init}}$, whose range is adjusted to 30–365. This modification is necessary due to the reduced amount of training data available under the revised data split. For the same reason, the window size of the benchmark model is reduced from two years to one year.
For the new test dataset, which spans a three-year period, RMSE and MAE are evaluated separately for the pre-crisis, crisis (spanning the period from 2021-09-01 to 2023-01-15), and post-crisis subperiods, as presented in Table 3.
MAE RMSE
Model Pre-crisis Crisis Post-crisis Pre-crisis Crisis Post-crisis
MLP with RLin 6.19 24.81 12.27 9.43 34.52 18.08
MLP with RLin (BOA) 5.96 24.27 11.45 8.98 33.58 17.32
LEAR 7.88 32.54 17.01 11.78 44.73 25.08
DNN 6.51 28.78 17.89 10.09 40.42 25.39
Table 3: MAE and RMSE for the three evaluation periods in the German–Luxembourg market. Pre-crisis period spans 2021-01-16 to 2021-08-31, Crisis period covers 2021-09-01 to 2023-01-15, and Post-crisis period corresponds to 2023-01-16 to 2024-01-15. For each metric and period, the cell colors range from red (indicating poorer performance) to blue (indicating better performance). The best performing models are highlighted in bold.
The main finding reported in Table 3 is that, although the forecasting accuracy of the proposed models deteriorates during the crisis period, their performance remains substantially superior to that of the benchmark models, with approximately a 16% reduction in MAE and 17% reduction in RMSE.
@@ 6 Conclusion
We adopt an ensemble method, following the fully adaptive BOA approach, to aggregate individual forecast models and produce more accurate predictions. Moreover, we propose a novel partial online learning method that substantially reduces computational cost by combining warm starting with window-specific hyperparameters.
The models’ performance is evaluated against two state-of-the-art benchmarks: LEAR and DNN models, using German-Luxembourg and Spanish bidding zones data, which are among the largest European markets. The measures of accuracy demonstrate the superiority of all the proposed models over the benchmark models with a 11% reduction in RMSE and 17% decrease in MAE for the German-Luxembourg market and a decline of 12% in RMSE and 14% in MAE for Spain. The results of the Diebold Mariano test are consistent with measures of accuracy metrics. Both LEAR and DNN fail to outperform any of the proposed models. Moreover, the superiority of the proposed models over the benchmarks remains evident even during the most recent European energy crisis.
In addition to their strong predictive accuracy, the suggested models achieve an outstanding runtime. It took the fast models only a few seconds to fit and generate the forecast, while even the slowest and most accurate models complete training and predictions within a few minutes. By contrast, the runtimes of benchmark models are substantially higher, such that each one of them took a couple of hours. This suggests that the proposed partial online learning approach was very effective in reducing computational time.
@@ Acknowledgements
This research was partially funded in the course of TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (520388526) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)
@@
Rangeinit Rangeup Applies to
lr{lr} $10^{-5},\,10^{-2}] [$$10^{-4},\,10^{-2}]$ All models
$D$ [30, 730][30,\,730] [1, 365][1,\,365] All models
$\lambda_{\texttt{1}}$ $10^{-5},\,10^{-2}] [$$10^{-4},\,10^{-2}]$ All models
$\lambda_{\texttt{2}}$ $10^{-5},\,10^{-2}] [$$10^{-4},\,10^{-2}]$ All models
$\alpha$ [0, 2][0,\,2] [0, 2][0,\,2] Only OLS-based models
nrnr [1, 128][1,\,128] [1, 128][1,\,128] Only models with MLP
Table 4: Hyperparameter search space used in the Optuna optimization. The parameters lrlr, $D$, $\lambda_{1}$, and $\lambda_{2}$ denote, respectively, the learning rate, the length of the training window (in days), and the penalty terms in the loss function. The parameter $\alpha$ controls the contribution of the OLS-based initialization and is only used for models with OLS initialization. Finally, nrnr denotes the number of neurons in the hidden layer. Rangeinit corresponds to the tuning range of the initial window, while Rangeup represents the tuning range of the updated windows.
@@
Figure 12: Pairwise DM Test P-Values Across Forecasting Models for Spanish market using L1 loss function
Figure 13: Pairwise DM Test P-Values Across Forecasting Models for the German-Luxembourg market using L2 loss function
Figure 14: Pairwise DM Test P-Values Across Forecasting Models for Spanish market using L2 loss function